# Prevention is better than cure? Feedback from high specific energy winds in cosmological simulations with Arkenstone

### v1

@@ 2.2 Galaxy formation model
Finally, we emphasise that the parameters of the TNG model (with the inclusion of its wind model) were calibrated to reproduce a set of observables, namely the SFR density as a function of $z$, the galaxy stellar mass function (SMF), the stellar mass to halo mass relation (SMHM), the BH mass to stellar mass relation, halo gas fractions, and galaxy stellar sizes. As of yet, we have not undertaken a similar calibration with Arkenstone. In this paper we show a set of numerical experiments showing the impact of changes in mass and energy loadings. A full flagship cosmological simulation, with parameters either calibrated or informed by idealised high-resolution ISM simulations, will be presented in a future work.
@@ 2.3 Arkenstone
Arkenstone is a framework for launching multiphase winds in cosmological simulations. It uses hydrodynamically decoupled wind particles to source a wind just outside the ISM, as the complex interaction between winds and the ISM is largely unresolved in coarse resolution cosmological simulations (as described in Section 1). The use of hydrodynamically decoupled wind particles has the added benefit of allowing us to control the properties of galactic winds more closely, so we can more easily investigate the impact of these properties on the CGM. We note that Arkenstone uses a completely independent wind particle implementation to that used in TNG, and is fully compatible with all other TNG physics.
@@
where $\dot{m}_{\star}$ is the star formation rate (SFR) of the cell, and $\eta_{\mathrm{M}}$ is the mass loading factor, which is a free parameter that we test in this work. This rate is then stochastically sampled to produce wind particles.
Upon creation, wind particles are given a “launch” energy in addition to the conserved quantities it receives from its parent cell. In contrast to S24, we launch wind particles isotropically, like TNG, with each new particle being given a kick in a random direction. The launch energies correspond to being given a velocity kick with magnitude
@@
where $\eta_{\mathrm{E,kin}}$ and $\eta_{\mathrm{E,th}}$ are the kinetic and thermal energy loadings, respectively, that are free parameters we vary in this study. Throughout this paper, we assume that the Mach number of the wind when it is launched is fixed at a value $\mathcal{M}=1$, which fixes the ratio $\eta_{\mathrm{E,kin}}/\eta_{\mathrm{E,th}}=5/9$. Like S24, we adopt $u_{\star}=5.26\times 10^{5}\,\mathrm{(km\,s^{-1})^{2}}$ as the characteristic specific energy associated with stellar feedback (dominated by supernovae), which corresponds to one supernova of $10^{51}$ erg per 95.5 M stellar mass formed (consistent with the value used in Kim et al., 2020a). Unlike EAGLE and TNG, we do not include a metallicity dependence of the energy loading (see Schaye et al., 2015; Pillepich et al., 2018a, for further details of their methods). Additionally, in this first work we take the metallicity of wind particles to be equal to that of the ISM. In reality, the hot, supernova driven phase is expected to have an enhanced metal fraction compared to the ISM (e.g. Li & Bryan, 2020; Kim et al., 2020b; Steinwandel et al., 2024a), but we postpone a detailed study of this aspect to future work.
@@
In Arkenstone, wind particles are created at a higher mass resolution (i.e. lower mass) than the base resolution of the simulation. This allows for better sampling of the energy injection rate in high specific energy winds, as well as forming part of the cosmological Arkenstone-Hot refinement scheme described below. A key result of S24 is that to accurately follow the propagation of a high specific energy wind, a simulation must have sufficient spatial resolution through the sonic point near the base of the wind (defined to be where the outflow velocity passes the radially decreasing sound speed as it adiabatically expands), where material is rapidly accelerated away from the galaxy. Arkenstone alleviates some of the difficulties in simulating high specific energy winds by ensuring the spatial resolution in the hot, low density phase of the winds remains a small fraction of the distance to the galaxy, as discussed in Sections 1 and 2.4, thus ensuring we accurately capture the wind’s acceleration. In this work, as in S24, all Arkenstone wind particles have a mass $100$ times smaller than the base resolution of the simulation (and of the wind particles in the TNG model). We note that the ratio of wind particle mass to base resolution mass was extensively tested in S24, who showed the mass and energy loadings measured in idealised simulations were very similar as this input parameter was varied.
@@
After launch, wind particles are recoupled when their host cell has a number density below 10 per cent of the star formation density threshold ($0.1n_{\rm H,0}\approx 10^{-2}\,\mathrm{cm}^{-3}$). We note this is higher than TNG’s threshold of 5 per cent of $n_{\rm H,0}$.
@@ 2.4 Recoupling and refinement
Table 1: List of simulation runs presented in this paper, with their wind mass loading, energy loading, and launch velocity. TNG scalings are calculated at $z=0$ as TNG wind launch velocities have a redshift dependence (Pillepich et al., 2018a); Arkenstone values/scalings apply at all redshifts. Note: the scalings with halo mass for TNG and the variable energy loading Arkenstone runs are inferred and will actually be scattered about this relation, as both actually scale with the locally measured DM velocity dispersion (and metallicity in the case of TNG).
Run $\eta_{\mathrm{M}}$ $\eta_{\mathrm{E}}$ $v_{\mathrm{launch}}$ [km s-1]
TNG $10.2(M_{\rm h}/10^{11}\,\mathrm{M}_{\odot})^{-5/6}$ (see Fig. 1) max $\left[350,378(M_{\rm h}/10^{11}\,\mathrm{M}_{\odot})^{1/3}\right]$
Constant $\eta_{\rm E}$ values (Oranges) 0.3 0.1 1061
0.3 0.3 613
0.3 0.9 354
Constant $\eta_{\rm M}$ values (Pink/Purple) 0.1 0.3 354
0.3 0.3 613
0.9 0.3 1061
Varying $\eta_{\rm E}$ values (Blues) 0.3 $0.4309(M_{\rm h}/10^{11}\,\mathrm{M}_{\odot})^{-1/6}$ 734$(M_{\rm h}/10^{11}\,\mathrm{M}_{\odot})^{-1/12}$
0.3 $0.4309(M_{\rm h}/10^{11}\,\mathrm{M}_{\odot})^{-1/3}$ 734$(M_{\rm h}/10^{11}\,\mathrm{M}_{\odot})^{-1/6}$
0.3 $0.4309(M_{\rm h}/10^{11}\,\mathrm{M}_{\odot})^{-1/2}$ 734$(M_{\rm h}/10^{11}\,\mathrm{M}_{\odot})^{-1/4}$
@@
The key component of the Arkenstone-Hot model described in S24 is the introduction of displacement recoupling. This prevents the dilution of the high specific energy of a wind particle, by displacing the existing contents of the host cell to its neighbouring cells before a wind particle recouples. Importantly, this also maintains the existing regular mesh structure so the hydrodynamics scheme continues to work correctly, which would not be the case if the standard Arepo refinement was used. When recoupling occurs, a passive scalar “dye" is injected, which is then advected as a tracer with the hydrodynamics scheme as a conserved quantity. Cells are flagged as “hot wind cells" if the ratio of this quantity to the cell mass, $f_{\mathrm{w}}$, is greater than a threshold value $f_{\mathrm{w,thresh}}$. In this work this threshold is set as $f_{\mathrm{w,thresh}}=1000$, the choice of which is described below and was extensively tested. Displacement recoupling only occurs [^4] Additional criteria apply that can prevent a cell from displacing material to its neighbours, though these occur rarely. See S24 for details. for host cells that do not already have a high temperature, or those flagged as part of the hot wind. Otherwise, a “simple” recoupling occurs, whereby wind particle quantities are directly injected into its host cell.
@@
However, we do not want to maintain higher resolution in the simulation indefinitely and refine the entire CGM, as this becomes prohibitively expensive. This requires a method of relaxing the refinement as the wind gets further from the galaxy. The fiducial Arkenstone implementation was designed and tested in isolated galaxy simulations that had an easily defined frame of reference, making it easy to implement a radial (i.e. galactocentric distance) criterion to prevent lower mass cells from de-refining. However, for the use of the Arkenstone model in cosmological simulations, a different method of managing the enhanced resolution is needed, as each galaxy has its own frame of reference. The simplest option could be to use a fixed refinement radius for all galaxies, though given haloes have different sizes this would mean maintaining high resolution for large distances around low mass galaxies, to ensure the same radius also encompasses the sonic point of massive haloes. To do this more efficiently for haloes of different masses therefore requires the wind particles to have knowledge of the halo they are created in.
Our new refinement scheme still makes use of the passive scalar dye injected by wind particles, described earlier, but with the addition of the value decaying over time. The scheme is designed to resolve the base of the galactic wind around galaxies of different masses, and without refining the entire CGM. To do this, we conservatively aim to maintain high resolution in the wind until $0.1R_{\mathrm{200}}$ [^5] We define $R_{\mathrm{200}}$ as the radius within which the mean density is 200 times the critical density of the universe at a given redshift. around a given galaxy. The particular choice of $0.1R_{\mathrm{200}}$ is somewhat arbitrary, as there is no physical reason for maintaining the resolution to this distance, but the motivations for it are two-fold. Firstly, we cannot easily predict the location of the sonic point around a given galaxy in a cosmological simulation, as it is affected by the galaxy’s geometry and environment. Secondly, the virial radius scales with halo mass, meaning we avoid the issues raised by using a fixed radius. Additionally, an estimate of the virial radius is available from local properties of gas cells that launch wind particles, as described below.
In post-processing of an initial simulation, we verified that the radius of $0.1R_{\mathrm{200}}$ comfortably encompasses a hot wind’s sonic point across a wide range of galaxies. There may be more efficient ways of determining a refinement radius, but the current scheme empirically works well without significant additional cost to the simulation.
@@
where $t_{0}$ is a decay constant we set to a fixed $t_{0}=10$ Myr. This therefore requires us to change the initial value of the dye, so that by the time the wind reaches $\sim 0.1R_{\mathrm{200}}$ the scalar tracer falls below a value of $f_{\rm w,thresh}=1000$ and cells are no longer flagged. We emphasise that the specific values of $t_{0}$ and $f_{\rm w,thresh}$ do not matter, it is the combination of them with the current value of the dye that determines a time until a cell begins to derefine. Though we note that using high values of $f_{\rm w,thresh}$ (together with high values of $f_{\rm w}$ at recoupling; see below) is preferred so that mixing with pristine, unflagged gas (which can also decrease the dye value) does not significantly affect the timer.
To estimate $R_{\mathrm{200}}$ from local properties at launch, we use the 1D dark matter velocity dispersion, which has been found to be correlated with the maximum circular velocity of a halo ($v_{\mathrm{max}}\simeq 1.45\sigma_{\mathrm{DM,1D}}$, Okamoto et al., 2010), and been used in a number of simulations for the launching of winds themselves (e.g. Oppenheimer & Davé, 2008; Vogelsberger et al., 2013). It can be shown for a Navarro-Frenk-White (NFW) halo (Navarro et al., 1997) that the maximum circular velocity $v_{\mathrm{max}}$ can be linked to the virial velocity $v_{\mathrm{200}}$ via
@@
During the simulation we will not know the concentration parameter $c$ for a given halo a priori, so we currently use a fixed value of $c=5$. Again we emphasise that the specific values entering the calculation do not matter, only the approximate scaling, and we have empirically found this to work.
@@
$v_{\mathrm{B}}=\sqrt{v_{\mathrm{w}}^{2}+2\gamma u_{\mathrm{w}}},$ (9)
which is the characteristic velocity for a wind launched with these properties, and $\gamma=5/3$ is the adiabatic index of the gas. The initial tracer value given to the wind particles to pass on to their host cell when recoupling occurs is therefore set to be
$f_{\mathrm{w,init}}=f_{\mathrm{w,thresh}}e^{\frac{0.1R_{\mathrm{200}}}{v_{\mathrm{B}}t_{0}}}.$ (10)
Gas thus travels approximately $0.1R_{200}$ before the tracer decays below $f_{\rm w}$, so cells maintain a higher resolution within that region. If the value of the tracer carried by a recoupling wind particle is higher than the existing tracer in its host cell, the tracer is overwritten.
@@
2.5 TNG winds
We now summarise the key parts of the TNG wind model (fully described in Pillepich et al., 2018a). As with Arkenstone, the TNG wind model also uses hydrodynamically decoupled wind particles (using completely independent implentations), launched isotropically from star forming gas cells. The TNG wind model does not include any of the features that Arkenstone uses for resolving high specific energy winds (higher resolution wind particles, displacement recoupling, wind refinement etc.).
Like Arkenstone, TNG uses a density threshold for recoupling (5% of the star formation threshold, compared to Arkenstone’s 10%), at which point the wind particle dumps all of its mass, energy, momentum and metals into its host cell (referred to as simple recoupling in the Arkenstone scheme).
Unlike in Arkenstone, the wind particles in TNG radiatively cool while hydro-decoupled, with their effective density adopted from the nearest gas cell. We omit this in Arkenstone for three reasons. Firstly, for high specific energy winds (which Arkenstone-Hot is designed to treat), the density of the local ISM, as measurable in the simulation, is not relevant – we are implicitly assuming that the hot, fast wind is propagating through unresolved low density channels and superbubble breakouts. Secondly, for these hot winds, the radiative cooling time is too long for this additional cooling to be important. Finally, our philosophy with Arkenstone is to avoid modelling the evolution of the wind within the ISM as much as possible, instead focusing on the properties of the wind that we source outside of the ISM; radiative cooling while in the ISM could be folded into the input energy loadings if desired.
The mass loading is then set as,
$\eta_{\rm M}=\frac{2}{v_{\rm w,TNG}^{2}}\left(1-\tau_{\rm w}\right)\eta_{\rm E}u_{\star},$ (13)
where $\tau_{\rm w}=0.1$ is the thermal fraction of the wind energy. Note that here we have expressed Equation 13 in a similar parametrisation to our Equations 2 and 3. The combination $u_{\star}\eta_{\rm E}$ is equivalent to the quantity $e_{w}$ in equation 3 of Pillepich et al. (2018a), which has a metallicity dependence (higher wind energy at lower metallicity). Pillepich et al. (2018a) also assumes a SN energy per unit stellar mass formed that is 13% higher than our reference $u_{\star}$ (relating to minor differences in IMF cutoff masses). Since $u_{\star}$ and $\eta_{\rm E}$ are degenerate as far as the actual wind energy input is concerned (and the difference is slight), for the purpose of comparing energy loadings between Arkenstone and TNG on an even footing, we fix $u_{\star}$ to our value and absorb the difference into the energy loading (along with the metallicity dependence).

### v2

@@ 2.3 TNG winds
Finally, we emphasise that the parameters of the TNG model were calibrated to reproduce a set of observables, namely the SFR density as a function of $z$, the galaxy stellar mass function (SMF), the stellar mass to halo mass relation (SMHM), the BH mass to stellar mass relation, halo gas fractions, and galaxy stellar sizes (Pillepich et al., 2018a). As of yet, we have not undertaken a similar calibration with Arkenstone. In this paper we show a set of numerical experiments showing the impact of changes in mass and energy loadings. A full flagship cosmological simulation, with parameters either calibrated or informed by idealised high-resolution ISM simulations, will be presented in a future work.
@@ 2.4 Arkenstone
Arkenstone is a framework for launching multiphase winds in cosmological simulations. It also uses hydrodynamically decoupled wind particles to source a wind just outside the ISM, as the complex interaction between winds and the ISM is largely unresolved in coarse resolution cosmological simulations (as described in Section 1). The use of hydrodynamically decoupled wind particles has the added benefit of allowing us to control the properties of galactic winds more closely, so we can more easily investigate the impact of these properties on the CGM. We note that Arkenstone uses a completely independent wind particle implementation to that used in TNG, and is fully compatible with all other TNG physics.
@@
where $\dot{m}_{\star}$ is the star formation rate (SFR) of the cell, and $\eta_{\mathrm{M}}$ is the mass loading factor, which is a free parameter that we test in this work. This rate is then stochastically sampled to determine how many wind particles to launch.
Upon creation, each wind particle is given a “launch” energy in addition to the conserved quantities it receives from its parent cell. In contrast to S24, we launch wind particles isotropically, like in TNG, with each new particle given a kick in a random direction. We note that these velocity kicks are actually added onto the velocity of the birth cell of the wind particle in both TNG and Arkenstone. The launch energies correspond to a velocity kick with magnitude
@@
where $\eta_{\mathrm{E,kin}}$ and $\eta_{\mathrm{E,th}}$ are the kinetic and thermal energy loadings, respectively. These are treated as free parameters that we vary in this study. Throughout this paper, we assume that the Mach number of the wind at launch is fixed at a value $\mathcal{M}=1$, which fixes the ratio $\eta_{\mathrm{E,kin}}/\eta_{\mathrm{E,th}}=5/9$. Like S24, we adopt $u_{\star}=5.26\times 10^{5}\,\mathrm{(km\,s^{-1})^{2}}$ as the characteristic specific energy associated with stellar feedback (dominated by supernovae), which corresponds to one supernova of $10^{51}$ erg per 95.5 M stellar mass formed (consistent with the value used in Kim et al., 2020a). Unlike EAGLE and TNG, we do not include a metallicity dependence of the energy loading (see Schaye et al., 2015; Crain et al., 2015; Pillepich et al., 2018a, for further details of their methods).
Additionally, in this first work we take the metallicity of wind particles to be equal to that of the ISM cell from which each particle is launched. In reality, the hot, supernova driven wind phase is expected to have an enhanced metal fraction compared to the ISM (e.g. Li & Bryan, 2020; Kim et al., 2020b; Steinwandel et al., 2024a). In Appendix A we perform a numerical experiment in which we increase the metallicity of launched wind particles by a constant (and somewhat arbitrary) factor of 5 and look at its effect on galaxy metallicities. In that Appendix, we show how this metal enhancement can change the normalisation of the mass-metallicity relation in Arkenstone. However, such a constant factor is not fully consistent with the model; unfortunately, simulating metal outflows as sourced directly by the SNe instead of being proportional to the ISM metallicity is technically challenging in our current implementation and we therefore postpone a detailed study of metallicity to future work.
@@
In Arkenstone, wind particles are created at a higher mass resolution (i.e. lower mass) than the base resolution of the simulation. This allows for better sampling of the energy injection rate in high specific energy winds, as well as forming part of the cosmological Arkenstone-Hot refinement scheme described below. A key result of S24 is that to accurately follow the propagation of a high specific energy wind, a simulation must have sufficient spatial resolution through the sonic point near the base of the wind (defined to be where the outflow velocity passes the radially decreasing sound speed as it adiabatically expands), where material is rapidly accelerated away from the galaxy. Arkenstone alleviates some of the difficulties in simulating high specific energy winds by ensuring the spatial resolution in the hot, low density phase of the winds remains a small fraction of the distance to the galaxy, as discussed in Sections 1 and 2.5, thus ensuring we accurately capture the wind’s acceleration.
In this work, as in S24, all Arkenstone wind particles have a mass $100$ times smaller than the base resolution of the simulation (and of the wind particles in the TNG model). During extensive testing we have found a slight dependence of quantities like galaxy stellar masses with the value of this ratio in cosmological simulations, which was not present in the isolated galaxies of S24. This lack of convergence is potentially a consequence of enhanced cooling at higher resolution, and is something we are investigating further. We note, however, that the changes due to this are much smaller than those arising from changes in wind loading quantities, so for this work we keep all simulations with a ratio of wind resolution to base resolution of 100.
@@
After launch, wind particles travel (without interacting hydrodynamically) until the cell they are in has a number density below 10 per cent of the star formation density threshold - $0.1n_{\rm H,0}\approx 10^{-2}\,\mathrm{cm}^{-3}$ (as in Springel & Hernquist, 2003; Smith et al., 2024a). We note this is slightly higher than TNG’s threshold of 5 per cent of $n_{\rm H,0}$. At this point, the wind particle returns its conserved quantities to the host cell via one of the recoupling methods described in the following section.
@@ 2.5 Recoupling and refinement
Table 1: List of simulation runs presented in this paper, with their wind mass loading, energy loading, and launch velocity kick. Halo mass scalings for variable loadings runs are calculated at $z=0$; the Arkenstone scalings with $\sigma_{\mathrm{DM,1D}}$ apply at all redshifts. The scalings with halo mass for TNG and the variable energy loading Arkenstone runs are inferred and will actually be scattered about this relation, as both actually scale with the locally measured DM velocity dispersion (and metallicity in the case of TNG).
Run $\eta_{\mathrm{M}}$ $\eta_{\mathrm{E}}$ $\Delta v_{\mathrm{launch}}$ [km s-1]
TNG $10.2\left(\frac{M_{\rm h}}{10^{11}\,\mathrm{M}_{\odot}}\right)^{-5/6}$ (see Fig. 1) max $\left[350,378\left(\frac{M_{\rm h}}{10^{11}\,\mathrm{M}_{\odot}}\right)^{1/3}\right]$
Fixed $\eta_{\rm E}$ values, 0.3 0.1 354
Constant $\eta_{\rm M}$, 0.3 0.3 613
(Oranges) 0.3 0.9 1061
Fixed $\eta_{\rm M}$ values, 0.1 0.3 1061
Constant $\eta_{\rm E}$, 0.3 0.3 613
(Pink/Purple) 0.9 0.3 354
Varying $\eta_{\rm E}$ values, 0.3 $0.4309\left(\frac{\sigma_{\rm DM,1D}}{\sigma_{\rm DM,1D,0}}\right)^{-1/2}\propto M_{\rm h}^{-1/6}$ $734\left(\frac{\sigma_{\rm DM,1D}}{\sigma_{\rm DM,1D,0}}\right)^{-1/4}\propto M_{\rm h}^{-1/12}$
Constant $\eta_{\rm M}$, 0.3 $0.4309\left(\frac{\sigma_{\rm DM,1D}}{\sigma_{\rm DM,1D,0}}\right)^{-1\,\,\,\,\,\,}\propto M_{\rm h}^{-1/3}$ $734\left(\frac{\sigma_{\rm DM,1D}}{\sigma_{\rm DM,1D,0}}\right)^{-1/2}\propto M_{\rm h}^{-1/6\,\,\,}$
(Blues) 0.3 $0.4309\left(\frac{\sigma_{\rm DM,1D}}{\sigma_{\rm DM,1D,0}}\right)^{-3/2}\propto M_{\rm h}^{-1/2}$ $734\left(\frac{\sigma_{\rm DM,1D}}{\sigma_{\rm DM,1D,0}}\right)^{-3/4}\propto M_{\rm h}^{-1/4\,\,\,}$
A key component of the Arkenstone-Hot model described in S24 is the introduction of the displacement recoupling method. When the host cell of a wind particle meets the density criterion described in Section 2.4, the wind particle injects its mass, momentum, energy, and metals into that cell. However, if a high resolution, high specific energy wind particle were to inject its energy immediately into a more massive host cell, it would be diluted, leading to significant numerical overcooling. To prevent dilution of the high specific energy of a wind particle, we first displace the existing contents of the host cell to its neighbouring cells before the wind particle recouples. Importantly, this also maintains the existing regular mesh structure so the hydrodynamics scheme continues to work correctly, which would not be the case if the standard Arepo refinement was used.
@@
When recoupling occurs, a passive scalar “dye" is injected, which is then advected as a tracer with the hydrodynamics scheme as a conserved quantity. Cells are flagged as “hot wind cells" if the ratio of this quantity to the cell mass, $f_{\mathrm{w}}$, is greater than a threshold value $f_{\mathrm{w,thresh}}$. In this work this threshold is set as $f_{\mathrm{w,thresh}}=1000$, the choice of which is described below. Displacement recoupling only occurs for host cells that do not already have a high temperature, or those already flagged as part of the hot wind. In S24, an additional requirement for cells to maintain a higher pressure than their surroundings (by retaining more mass during the displacement process) was used. However, we found in cosmological simulations that this occurred too frequently and had the effect of diluting our wind energy, so in this work we do not use this criterion. [^4] Additional criteria can also apply to prevent a cell from displacing material to its neighbour (see S24, for details), though these occur rarely. If a displacement recoupling does not happen a “simple” recoupling occurs instead, whereby wind particle quantities are directly injected into its host cell, as happens in TNG.
@@
However, we do not want to maintain higher resolution in the simulation indefinitely and refine the entire CGM, as this becomes prohibitively expensive. This requires a method of relaxing the refinement as the wind gets further from the galaxy. The fiducial Arkenstone implementation was designed and tested in isolated galaxy simulations that had an easily defined frame of reference, making it simple to implement a radial (i.e. galactocentric distance) criterion to prevent lower mass cells from de-refining. However, for the use of the Arkenstone model in cosmological simulations, a different method of managing the enhanced resolution is needed, as each galaxy has its own frame of reference. The simplest option could be to use a fixed refinement radius for all galaxies, though because haloes have different sizes this would mean maintaining high resolution for large distances around low mass galaxies, to ensure the same radius also encompasses the sonic point of massive haloes. To do this more efficiently for haloes of different masses therefore requires the wind particles to have knowledge of the halo in which they are created.
Our new refinement scheme still makes use of the passive scalar dye injected by wind particles, described earlier, but instead the value decays over time. The scheme is designed to resolve the base of the galactic wind around galaxies of different masses, without refining the entire CGM. To do this, we conservatively aim to maintain high resolution in the wind until $0.2R_{\mathrm{200}}$ [^5] We define $R_{\mathrm{200}}$ as the radius within which the mean density is 200 times the critical density of the universe at a given redshift. around a given galaxy. The particular choice of $0.2R_{\mathrm{200}}$ is somewhat arbitrary, as there is no physical reason for maintaining the resolution to this distance, but the motivations for it are two-fold. Firstly, we cannot easily predict the location of the sonic point around a given galaxy in a cosmological simulation, as it is affected by the galaxy’s geometry and environment. Secondly, the virial radius scales with halo mass, so that we avoid the issues raised by using a fixed radius. Additionally, an estimate of the virial radius is available from local properties of gas cells that launch wind particles, as described below.
In post-processing of an initial simulation, we verified that the radius of $0.2R_{\mathrm{200}}$ comfortably encompasses the sonic point of the hot wind across a wide range of galaxy masses. There may be more efficient ways of determining a refinement radius that we will investigate in future work, but the current scheme works well empirically and so is used throughout this paper.
@@
where $t_{0}$ is a decay constant we set to a fixed $t_{0}=10$ Myr. This therefore requires us to change the initial value of the dye, so that by the time the wind reaches $\sim 0.2R_{\mathrm{200}}$ the scalar tracer falls below a value of $f_{\rm w,thresh}=1000$ and cells are no longer flagged. We emphasise that the specific values of $t_{0}$ and $f_{\rm w,thresh}$ do not matter, it is the combination of them with the current value of the dye that determines a time until a cell begins to derefine. Though we note that using high values of $f_{\rm w,thresh}$ (together with high values of $f_{\rm w}$ at recoupling; see below) is preferred so that mixing with pristine, unflagged gas (which can also decrease the dye value) does not significantly affect the timer.
To estimate $R_{\mathrm{200}}$ from local properties at launch, we use the 1D dark matter velocity dispersion, which has been found to be correlated with the maximum circular velocity of a halo ($v_{\mathrm{max}}\simeq 1.45\sigma_{\mathrm{DM,1D}}$, Okamoto et al., 2010), and been used in a number of simulations for the launching of winds themselves (e.g. Oppenheimer & Davé, 2008; Vogelsberger et al., 2013). This is calculated in a kernel-weighted fashion over the nearest 64 dark matter particles to the host gas cell. It can be shown for a Navarro-Frenk-White (NFW) halo (Navarro et al., 1997) that the maximum circular velocity $v_{\mathrm{max}}$ can be linked to the virial velocity $v_{\mathrm{200}}$ via
@@
During the simulation, we will not know the concentration parameter $c$ for a given halo a priori, so we currently use a fixed value of $c=2$, which minimises the function $f(c)$ and so provides the most conservative estimate of $R_{\mathrm{200}}$.
@@
$v_{\mathrm{B}}=\sqrt{(\Delta v_{\mathrm{w}})^{2}+2\gamma u_{\mathrm{w}}},$ (11)
which is the characteristic velocity for a wind launched with these properties, and $\gamma=5/3$ is the adiabatic index of the gas. In reality, due to a number of factors, the wind will not actually be at this velocity, and so we introduce an additional factor $k=0.25$ to encompass such uncertainties. Again, we emphasise that the specific values entering the calculation do not matter, only the approximate scaling, and we have empirically found this to work.
The initial tracer value given to the wind particles to pass on to their host cell when recoupling occurs is therefore set to be
$f_{\mathrm{w,init}}=f_{\mathrm{w,thresh}}\exp{\frac{0.2R_{\mathrm{200}}}{kv_{\mathrm{B}}t_{0}}}.$ (12)
Gas thus travels approximately $0.2R_{200}$ before the tracer decays below $f_{\rm w}$, so cells maintain a higher resolution within that region. If the value of the tracer carried by a recoupling wind particle is higher than the existing tracer in its host cell, the tracer is overwritten.